Write an equation of the line with the given properties. Your answer should be written in standard form.
m = −1/5 passing through P(−6, −9)

Answer by Cromlix(4381)

Using the form y - b = m(x - a)
y -(-9) = -1/5(x -(-6))
y + 9 = -1/5x - 6/5
y = -1/5x - 6/5 - 45/5
y = -1/5x - 51/5
or multiply through out by 5
5y = -x - 51.
Hope this helps.
:-)